Find four consecutive odd integers such that the product of the second and third exceeds the square of the first by 98.
Answer by mananth(16946)

let the integers be x,x+2,x+4,x+6
product of second and third = (x+2)(x+4)
Square of first = x^2
..
x^2+98 = (x+2)(x+4)
x^2+98 = x^2+6x+8
6x+8 = 98
6x =98-8
6x =90
x= 15
.....
15,17 19,21